Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Good old probability

Expert replies
by alex.gellatly » Sat Jun 23, 2012 8:10 pm
There are 10 women and 3 men in room A. One person is picked at random from room A and moved to room B, where there are already 3 women and 5 men. If a single person is then to be picked from room B, what is the probability that a woman will be picked?

13/21
49/117
40/117
15/52
5/18
Join the discussion
Source: — Problem Solving |

by gmat_and_me » Sat Jun 23, 2012 8:54 pm
The resultant probability has to be the product of probabilities
of selecting one woman from room A and afterwards selecting a
woman from room B.

Probability of selecting one woman from room A = 10/13
Resultant probability = (10/13) * (4/9) = 40/117

HTH
alex.gellatly wrote:There are 10 women and 3 men in room A. One person is picked at random from room A and moved to room B, where there are already 3 women and 5 men. If a single person is then to be picked from room B, what is the probability that a woman will be picked?

13/21
49/117
40/117
15/52
5/18
Join the discussion

by gmat_and_me » Sat Jun 23, 2012 9:04 pm
I think, I wrote it too fast. My bad. There is one more
possibility. You can select one man from room A and then
select a woman from room B. In that case P will be

(3/13) * (3/9) = 9/117

Since, either of these two can happen, the resultant probability
will then be the sum of the two => 40/117 + 9/117 = 49/117.

HTH
gmat_and_me wrote:The resultant probability has to be the product of probabilities
of selecting one woman from room A and afterwards selecting a
woman from room B.

Probability of selecting one woman from room A = 10/13
Resultant probability = (10/13) * (4/9) = 40/117

HTH
alex.gellatly wrote:There are 10 women and 3 men in room A. One person is picked at random from room A and moved to room B, where there are already 3 women and 5 men. If a single person is then to be picked from room B, what is the probability that a woman will be picked?

13/21
49/117
40/117
15/52
5/18
Join the discussion

by alex.gellatly » Sat Jun 23, 2012 9:45 pm
gmat_and_me wrote:I think, I wrote it too fast. My bad. There is one more
possibility. You can select one man from room A and then
select a woman from room B. In that case P will be

(3/13) * (3/9) = 9/117

Since, either of these two can happen, the resultant probability
will then be the sum of the two => 40/117 + 9/117 = 49/117.

HTH
gmat_and_me wrote:The resultant probability has to be the product of probabilities
of selecting one woman from room A and afterwards selecting a
woman from room B.

Probability of selecting one woman from room A = 10/13
Resultant probability = (10/13) * (4/9) = 40/117

HTH
alex.gellatly wrote:There are 10 women and 3 men in room A. One person is picked at random from room A and moved to room B, where there are already 3 women and 5 men. If a single person is then to be picked from room B, what is the probability that a woman will be picked?

13/21
49/117
40/117
15/52
5/18
Yes, the OA is B. I made your same mistake at first as well.
Join the discussion